A Bibliography of Publications of Jean-Marc Wagner

Jean-Marc Wagner Institut Montefiore B28, R75b

Sart Tilman B-4000 Liège, Belgium

Tel: +32 (0)4 366 27 93 (secr. +32 (0)4 366 26 91) FAX: +32 (0)4 366 29 89

E-mail: JM.Wagner@ulg.ac.be (Internet)

27 November 2011 Version 1.03

Abstract

This bibliography records publications of Jean-Marc Wagner.

Title word cross-reference

2 [NW01]. 3 [WN99].

-D [NCW01b].

180-degrees [NW01].

Acoustics [IEE01]. acquisition [NW01]. algorithm [WN00]. analytiques [Wag02]. Attenuation [WNC^{+}01c].

Band [NCW01b]. beam [WN01].

Center [IEE01]. City [IEE01]. Completeness [NCW01a]. Condition [NCW01a]. Conference [IEE01]. construction [WN99].

Convention [IEE01]. Correction [WNC^{+}01c].

D [WN99]. d’atténuation [Wag02]. degrees [NW01]. dimensional [WN01]. d’une [WN99].

émission [Wag02]. Equatorial [NCW01b]. Exact [WNC^{+}01c]. exponential [WN00].

Formula [NCW01a].

Hole [WNC^{+}01c].

image [WN99]. intégrales [WN99]. International [IEE01]. inversion [WN00].

Lake [IEE01]. linogramme [WN99].

May [IEE01]. Méthode [WN99]. Méthodes [Wag02]. monophotonique [Wag02].

Neumann [WNC01a]. octogonal [WN99].

Palace [IEE01]. parallel [WN01]. parallel-beam [WN01]. partir [WN99]. plans [WN99]. pour [WN99]. proceedings [IEE01]. Processing [IEE01]. Projections [NCW01a].

ray [WN00]. Rebinning [WNC^{+}01c]. Reconstruction [NCW01a]. Rotating [WNC^{+}01c]. RSH [WNC^{+}01c].

Salt [IEE01]. Semi [NCW01b]. series [WNC01a]. ses [WN99]. Signal [IEE01]. Slant [WNC^{+}01c]. Slant-Hole [WNC^{+}01c]. SPECT [NW01]. Speech [IEE01].

Three [WN01]. Three-dimensional [WN01]. tomographie [Wag02]. transform [WN00]. tridimensionnelle [Wag02]. TTR [WN00].

USA [IEE01]. using [WNC^{+}01c]. Utah [IEE01].

X [WN00]. X-ray [WN00].

References

IEEE:2001:IIC

[IEE01] IEEE, editor. 2001 IEEE International Conference on Acoustics, Speech, and Signal Processing: proceedings: 7–11 May, 2001, Salt Palace Convention Center, Salt Lake City, Utah, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7803-7041-4, 0-7803-7042-2 (microfiche). LCCN ???? Six volumes.

Noo:2001:IRE

[NCW01a] F. Noo, R. Clackdoyle, and J.- M. Wagner. 3D image reconstruction from exponential Xray projections: a completeness condition and an inversion formula. In IEEE Medical Imaging Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001.

URL http://www.ulg.ac.be/
telecom/publi/publications/
wagner/MIC2001FRED.pdf.

Noo:2001:IDE

[NCW01b] F. Noo, R. Clackdoyle, and J.- M. Wagner. Inversion of the 3-D exponential X-ray transform for a semi equatorial band. In 3D Meeting. Asilomar, USA, 2001.

URL http://www.ulg.ac.be/
telecom/publi/publications/
wagner/abstract_3D2001_Fred.
pdf.

Noo:2001:IRS

[NW01] F. Noo and J.-M. Wagner. Image reconstruction in 2D SPECT with 180-degrees acquisition. Inverse Problems, 17:1357–1371, 2001. CODEN INPEEY. ISSN 0266-5611.

URL http://www.ulg.ac.be/
telecom/publi/publications/
wagner/ERT180.pdf.

Wagner:2002:MAP

[Wag02] J.-M. Wagner. Méthodes analytiques pour la correction d’atténuation en tomographie tridimensionnelle par émission monophotonique. PhD thesis, University of Liège, Liège, Belgium, February 2002.

Wagner:1999:MDL

[WN99] J.-M. Wagner and F. Noo.

Méthode du linogramme octogonal pour la construction d’une image 3D à partir de ses intégrales de plans. In Gretsi,

1999. URL http://www.ulg.ac.
be/telecom/publi/publications/
image/gretsi99/gretsi99.html;
 http://www.ulg.ac.be/telecom/
publi/publications/gretsi/
gretsi99.pdf.

Wagner:2000:TAI

[WN00] J.-M. Wagner and F. Noo.

TTR algorithm for the inversion of the exponential X-ray transform. In IEEE Medical Imaging Conference, Lyon, France, October 2000. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000.

URL http://www.ulg.ac.be/
telecom/publi/publications/
wagner/lyon2000/index.html;
 http://www.ulg.ac.be/telecom/
publi/publications/wagner/
IEEE2000MIC.pdf.

Wagner:2001:TDI

[WN01] J.-M. Wagner and F. Noo. Threedimensional image reconstruction from parallel-beam projections. IEEE Transactions on Nuclear Sciences, 48:743– 749, June 2001. CODEN IRNSAM. ISSN 0018-9499 (print), 1558-1578 (electronic).

                                        URL http://www.ulg.ac.be/
                                        telecom/publi/publications/
                                        wagner/articleTNS.pdf.

Wagner:2001:IRE

[WNC01a] J.-M. Wagner, F. Noo, and R. Clackdoyle. 3D image reconstruction from exponential X-ray projections using Neumann series. In IEEE [IEE01]. ISBN 0-7803-7041-4, 0-7803-7042-2 (microfiche). LCCN ????

                                        URL http://www.ulg.ac.be/
                                        telecom/publi/publications/
                                        wagner/Neumann.pdf. Six vol-

umes.

Wagner:2001:EIE

[WNC01b] J.-M. Wagner, F. Noo, and R. Clackdoyle. Exact inversion of the exponential X-ray transform for RSH SPECT. In 3D Meeting. Asilomar, USA, 2001.

                                        URL http://www.ulg.ac.be/
                                        telecom/publi/publications/
                                        wagner/abstract_3D2001.pdf.

Wagner:2001:ACR

[WNC^{+}01c] J.-M. Wagner, F. Noo, R. Clackdoyle, G. Bal, and P. Christian. Attenuation correction for rotating slant-hole (RSH) SPECT using exact rebinning. In IEEE Medical Imaging Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001.

                                        URL http://www.ulg.ac.be/
                                        telecom/publi/publications/
                                        wagner/MIC2001JM.pdf.